# Special Hamiltonian $S^{1}$-actions on symplectic 4-manifolds

### v3

@@ 5 A special Hamiltonian $S^{1}$-action on $\mathcal{W}_{n}$
It is known that all Lagrangian spheres in the cotangent bundle $T^{*}S^{2}$ are hamiltonian isotopic [7], so are those in $\mathcal{W}_{1}$. Recall from Proposition 3.1 that up to conjugation any semi-simple special Hamiltonian $S^{1}$-action on $\mathcal{W}_{1}$ are symplectically equivalent to the standard G-action when restricted to some open neighborhood $U$ of the Lagrangian sphere $S$. We may assume that $(x,y)\in U$ whenever $|y|:=\sqrt{y^{2}}<\delta$ for some $\delta>0$.
Let $\eta_{t}$ be the time $t$ map of the flow of $Y$ with $\eta_{0}=id_{\mathcal{W}_{1}}$. $\eta_{t}$ is G-invariant for all $t$. For $(x,y)\in\mathcal{W}_{1}$
@@
Let $\psi_{t}$ be the time $t$ map of the flow of $Y_{\rho}$. $\frac{d\psi_{t}}{dt}=Y_{\rho}(\psi_{t})$. $\phi_{0}=id_{\mathcal{W}_{1}}$. $\frac{d\psi_{t}}{dt}=0$ when restricted to $S$. $\psi_{t}$ is G-equivariant for all $t$. For $(x,y)\in\mathcal{W}_{1}$,
@@
As in the proof of Theorem 1.6(i) we may assume that $=\textsf{G}$ G when restricted to the Lagrangian sphere $S:x^{2}=1,y=0$ in $\mathcal{W}_{1}$, and consider the exponential map $\EuFrak{e}_{S}$ over $S$ relative to a $G$-invariant $\omega_{1}$-compatible Riemannian metric $g$ on $\mathcal{W}_{1}$:
@@
$G_{\theta}=\EuFrak{e}_{S}\circ\textsf{G}_{\theta}\circ(\EuFrak{e}_{S})^{-1},\quad\forall\theta.$
@@
$F_{r}^{-1}\circ G_{\theta}\circ F_{r}=\textsf{G}_{\theta}\quad\text{on $U_{r}:=\{(x,y)\in\mathcal{W}_{1}||y|^{2}\leq r\}$}.$
By allowing $r_{1}$ and $r_{0}$ to be arbitrarily large we conclude that any semi-free special Hamiltonian $S^{1}$-action on $(\mathcal{W}_{1},\omega_{1})$ is eventually symplectically equivariant to the standard linear action of G. This completes the proof of Case 2.
@@ 6 Non-simply connected cases
Now we consider the case that $(M,\omega)$ is a connected exact symplectic 4-manifold with $c_{1}(M,\omega)=0$ but not 1-connected. Assume that $(M,\omega)$ equips with a semi-free special Hamiltonian $S^{1}$-action which we denote as $G\subset\text{Ham}(M,\omega)$.
Since the Maslov condition applies only when nonconstant orbits of $G$ are homologically trivial, we assume that $Fix(G)$ is nonempty and finite to accommodate the Maslov condition . Let
$k=n+1:=|Fix(G)|\geq 1$
denote the number of fixed points of the $G$-action, $h:M\to\mathbb{R}$ the moment map associated to $G$, and $X=X_{h}$ the Hamiltonian vector field of $h$ defined by $\omega(X,\cdot)=-dh$.
Fix any $\lambda\in\Omega^{1}(M)$ with $d\lambda=\omega$. As in the 1-connected case we may assume that $\lambda$ is $G$-invariant and $Fix(G)\subset h^{-1}(0)$, then $\lambda$ satisfies $\lambda(X)=h$. Let
$M_{red}:=h^{-1}(0)/G\cong h^{-1}(c)/G$
denote the symplectic reduction of $(M,\omega,G)$ via the reduction map
$\EuFrak{r}:M\to M_{red}$
associated to the $G$-action on $M$. $M_{red}$ is topologically an oriented surface of genus $g$ with $k$ marked points ($\EuFrak{r}(Fix(G))$) and $b\geq 1$ connected boundary components at infinity. Since $G$ acts freely on $h^{-1}(c)$ for all $c\neq 0$, $M_{red}$ is not 1-connected and hence
@@
Let $B\subset M_{red}$ be a bounded open domain of $M_{red}$ which is homotopic to $M_{red}$ and contains $\EuFrak{r}(Fix(G))$ in its interior. Up to homotopy equivalence we may identify $M$ with the corresponding $G$-invariant subdomain $\EuFrak{r}^{-1}(B)\cap h^{-1}((-r,r))$ of $M$ for some $r>0$.
Part 1: Handle decomposition of $B$ and Homology of $M$.  For each of the $k=1+n$ fixed points $p_{i}$, $0\leq i\leq n$, of the $G$-action on $M$, denote by $q_{i}:=\EuFrak{r}(p_{i})\in B$ the image of $p_{i}$ under the symplectic reduction map $\EuFrak{r}:M\to M_{red}$. In $B$ pick $n+1$ open discs $D_{i}$ centered at $q_{i}$, $i=0,1,\dots,n$, such that the closures of $D_{i}$’s are pairwise disjoint closed discs in $B$. As $B$ can be constructed by attaching 2-dimensional 1-handles to 2-dimensional 0-handles, we may view $D_{i}$ as the set of 2-dimensional 0-handles of $B$, and with $q_{i}$ as the core point of $D_{i}$. Then $B$ can be obtained by attaching the following two types of 1-handles to $\cup_{i}D_{i}$:
@@
    Let $\tau_{i}\subset B\setminus D_{0}$, $i=1,\dots,n$, be a set of pairwise disjoint embedded arcs with endpoints such that $\tau_{i}$ is disjoint from $q_{l}$ if $l\neq i$, $q_{i}=\partial_{+}\tau_{i}$ is one endpoint of $\tau_{i}$, and $\tau_{i}$ intersects transversally with $\partial D_{0}$ at the other endpoint $\partial_{-}\tau_{i}$ of $\tau_{i}$. Thicken each of $\tau_{i}$ a bit to get $n$ mutually disjoint 2-dimensional 1-handles $H_{\tau_{i}}$ with $\tau_{i}$ as the core curve of $H_{\tau_{i}}$, $H_{\tau_{i}}\cap D_{l}=\emptyset$ if $l\neq i$ or 00, and $H_{\tau_{i}}\cap D_{0}$ is the end interval of $H_{\tau_{i}}$ which contains the endpoint $\partial_{-}\tau_{i}$ of $\tau_{i}$. WLOG we may assume that that $D_{i}\cup H_{\tau_{i}}$ is 1-connected for $i=1,2,\dots,n$.
@@
    $D$ is diffeomorphic to a 2-disc. Note that $Fix(G)\subset\EuFrak{r}^{-1}(D)$, and $\EuFrak{r}^{-1}(D)$ can be identified with a 1-connected Stein domain $W_{n}\subset\mathcal{W}_{n}$ is a Stein domain.
@@
    A disjoint union $(2g+b-1)$ 1-handles with both ends attached to $\partial D_{0}$, missing all of the handles $H_{\tau_{i}}$ in (i):
    Denote these 1-handles as $H_{C_{j}}$ for $j=1,2,\dots,2g+b-1$, where $C_{j}$ is the core arc of $H_{C_{j}}$. Denote the boundary of $H_{C_{j}}$ as $\partial H_{C_{j}}=\partial_{+}H_{C_{j}}\cup\partial_{-}H_{C_{j}}$, $H_{C_{j}}$ is attached to $\partial D$ along $\partial_{-}H_{C_{j}}$, where $\partial_{-}H_{C_{j}}$ is the disjoint union of two short intervals containing $\partial C_{j}$. By handle sliding along $\partial D$ (and thinning $\tilde{H}_{\tau_{i}}$ and $H_{C_{j}}$ if necessary) we may assume that all $\partial_{-}H_{C_{j}}$ are attached to $\partial D_{0}$ and not touching any part of $\partial H_{\tau_{i}}$.
@@
Recall $\EuFrak{r}_{0}:h^{-1}(0)\to h^{-1}(0)/G=B$ the standard projection from $h^{-1}(0)\subset M$ to the reduced space. Fix a $G$-invariant $\omega$-compatible Riemannian metric on $M$ and let $\nabla h$ denote the gradient vector field of the moment map $h$ with respect to the metric. $M$ deformation retracts to the hypersurface $h^{-1}(0)$ via flows of gradients ±$\pm\nabla h$ of $h$, therefore $H_{i}(M,\mathbb{Z})\cong H_{i}(h^{-1}(0),\mathbb{Z})$ for all $i$.
Now $h^{-1}(0)=\EuFrak{r}_{0}^{-1}(B)$. As $G$ acts freely on $\EuFrak{r}_{0}^{-1}(B)\setminus Fix(G)$, we have the following results.
@@
    $\EuFrak{r}_{0}^{-1}(\tau_{i})\subset h^{-1}(0)$ is Lagrangian disc centered at $p_{i}$, $1\leq i\leq n$. If we extend $\tau_{i}$ to an embedded arc $\tilde{\tau}_{i}\subset D$ with $q_{i}$ and $q_{0}$ as endpoints, then $\tilde{S}_{i}:=\EuFrak{r}_{0}^{-1}(\tilde{\tau}_{i})$ is a Lagrangian 2-sphere with points $p_{i}$ and $p_{0}$ as poles for $i=1,2,\dots n$. These $n$ spheres are homologically independent in $\EuFrak{r}^{-1}_{0}(D)$ and hence in $W_{n}=\EuFrak{r}^{-1}_{G}(D)$. Indeed $W_{n}$ is homotopic to the bouquet $\cup_{i=1}^{n}\tilde{S}_{i}$ of $n$ 2-spheres.
@@
    $\EuFrak{r}_{0}^{-1}(C_{j})\subset h^{-1}(0)$ is Lagrangian cylinder with boundary attached to $\EuFrak{r}^{-1}_{0}(\partial D)$. If we extend $C_{j}$ to a simple closed curve $\tilde{C}_{j}\subset D\setminus\{q_{i}\mid i=0,\dots n\}$ then we get an embedded Lagrangian torus $\tilde{T}_{j}\subset h^{-1}(0)=\EuFrak{r}^{-1}_{0}(B)$ for $j=1,2,\dots,2g+b-1$. These curves $\tilde{C}_{j}$, $j=1,\dots,2g+b-1$, together represent a basis of the homology group $H_{1}(B,\mathbb{Z})\cong\mathbb{Z}^{2g+b-1}$. As $G$-orbits in $h^{-1}(0)$ contracts to points in $Fix(G)$, $H_{1}(h^{-1}(0),\mathbb{Z})\cong H_{1}(B,\mathbb{Z})$ and is generated by a lifting of the curves $\tilde{C}_{j}$ in $\tilde{T_{j}}$, $j=1,\dots,2g+b-1$. The tori $\tilde{T}_{j}$’s are also independent on $H_{2}(h^{-1}(0),\mathbb{Z})$. Also $\tilde{T_{j}}$’s and $\tilde{S}_{i}$’s are pairwise linearly independent in $H_{2}(h^{-1}(0),\mathbb{Z})$, and forming a basis for $H_{2}(h^{-1}(0),\mathbb{Z})\cong\mathbb{Z}^{2g+b-1+n}$. Therefore
    $H_{m}(M,\mathbb{Z})=\begin{cases}0&\text{ for }m>2,\\ \mathbb{Z}^{2g+b-1+n}\cong H_{2}(W_{n},\mathbb{Z})\oplus\mathbb{Z}^{2g+b-1}&\text{ for }m=2,\\ H_{1}(\Sigma=M_{red},\mathbb{Z})\cong\mathbb{Z}^{2g+b-1}&\text{ for }m=1.\end{cases}$
    In particular the homology groups of $(M,\omega,G)$ is completely determined by the triple $(k=n+1,g,b)$ with $k\geq 0$, $g\geq 0$, $b\geq 1$, and $g+b\geq 2$.
Parr 2: Stein structure on $(M,\omega=d\lambda,G)$
Below we study the existence of Stein structure on $(M,\omega=d\lambda,G)$. Recall the following theorem due to Eliashberg [2] (see also [3] Theorem 1.3 and [4] Theorem 2.3) about the existence of a Stein structure on a 4-manifold via handle attaching criteria, which can be stated as the following after Gompf ([4] Theorem 2.3):
@@
Below we will show that the handle decomposition of $B=D_{0}\cup(\cup_{i}\tilde{H}_{\tau_{i}})\cup(\cup_{j}H_{C_{j}})$ lifts to a corresponding handle decomposition of $M$ as a union of handles
@@
where $\textsf{H}_{0}=W_{0}=\EuFrak{r}^{-1}(D_{0})$ is a 0-handle, $\textsf{H}_{V_{i}}$ are 2-handles associated to $\tilde{H}_{\tau_{i}}$, and for each $j$, the pair $(\textsf{H}_{\gamma_{j}},\textsf{H}_{U_{j}})$, where $\textsf{H}_{\gamma_{j}}$ is a 1-handle and $\textsf{H}_{U_{j}}$ is a 2-handle, is associated to $H_{C_{j}}$. Moreover, all the 2-handles $\textsf{H}_{V_{i}}$ and $\textsf{H}_{U_{j}}$ of $M$ satisfy the contact framing condition, hence $M$ admits the structure of a Stein manifold.
Case 1: $\textsf{H}_{V_{i}}$.  Denote by $V_{i}:=\EuFrak{r}^{-1}_{0}(\tau_{i})$ the Lagrangian disc in $h^{-1}(0)$ with the boundary circle $\partial V_{i}$ attached to $\EuFrak{r}^{-1}_{0}(\partial D_{0})\subset\EuFrak{r}^{-1}_{G}(\partial D_{0})$. $V_{i}$ is the core disc of a 4-dimensional handle $\textsf{H}_{V_{i}}$ which can be identified with a subdomain of the cotangent bundle over $V_{i}$. We may assume that these handles are mutually disjoint when attached to $\EuFrak{r}^{-1}(\partial D_{0})$.
Pick any $V_{i}$ and denote it as $V$, also denote $\tau:=\tau_{i}=\EuFrak{r}_{0}(V)$, $p:=p_{i}$.
We may parametrize $V$ as
$V=\{(x_{1},x_{2})\mid x_{1}^{2}+x_{2}^{2}\leq 1\},\quad p=(0,0)\in Fix(G)$
@@
$\textsf{H}_{V}:=\EuFrak{r}^{-1}_{G}(H_{\tau})\cong V\times D^{2}_{y}\subset T^{*}L,\quad D^{2}_{y}=\{(y_{1},y_{2})\mid y_{1}^{2}+y_{2}^{2}<\epsilon^{2}\}$
@@
Identify $\textsf{H}_{V}$ as a subdomain of $\mathbb{C}^{2}$. Then on $\textsf{H}_{V}$ we may take $\omega=d\lambda$, with $\lambda=\frac{1}{2}\sum_{i=1}^{2}(x_{i}dy_{i}-y_{i}dx_{i})$, identify the $G$-action with the standard G-action on $\mathbb{C}^{2}$, and take the Euclidean metric as the $G$-invariant $\omega$-compatible Riemannian metric on $\textsf{H}_{V}$. Then
@@
Note that both $X$ and $\nabla h$ are tangent to $\EuFrak{r}^{-1}(\partial D_{0})$ and hence to $\partial_{-}\textsf{H}_{V}$. The 1-form $\lambda$ restricts to a contact 1-form $\alpha$ on $\partial V\times D^{2}_{\epsilon}\subset\partial W_{0}$. $\lambda(X)=0$ along $\partial V$ so $\partial V$ is a Legendrian curve to the contact form. As the outward normal vector field to $\EuFrak{r}^{-1}(D_{0})=W_{0}$ near $\partial V$ is $-x_{1}\partial_{x_{1}}-x_{2}\partial_{x_{2}}\subset\ker\lambda|_{\partial V}$, the contact plane field along $\partial V$ is spanned by the symplectic pair $\{\nabla h,X\}$. So the contact framing along $\partial V$ (oriented by $X$) is given by the vector field $\nabla h|_{\partial V}=-x_{1}\partial_{y_{2}}+x_{2}\partial_{y_{1}}$.

### v4

@@ 5 A special Hamiltonian $S^{1}$-action on $\mathcal{W}_{n}$
It is known that all Lagrangian spheres in the cotangent bundle $T^{*}S^{2}$ are hamiltonian isotopic [7], so are those in $\mathcal{W}_{1}$. Recall from proof of Theorem 1.10 (i) that up to conjugation any semi-free SHam1-action on $\mathcal{W}_{1}$ are symplectically equivalent to the standard $\Theta$-action when restricted to some open neighborhood $U$ of the Lagrangian sphere $S$. We may assume that $(x,y)\in U$ whenever $|y|:=\sqrt{y^{2}}<\delta$ for some $\delta>0$.
Let $\eta_{t}$ be the time $t$ map of the flow of $Y$ with $\eta_{0}=id_{\mathcal{W}_{1}}$. $\eta_{t}$ is $\Theta$-invariant for all $t$. For $(x,y)\in\mathcal{W}_{1}$
@@
Let $\psi_{t}$ be the time $t$ map of the flow of $Y_{\rho}$. $\frac{d\psi_{t}}{dt}=Y_{\rho}(\psi_{t})$. $\phi_{0}=id_{\mathcal{W}_{1}}$. $\frac{d\psi_{t}}{dt}=0$ when restricted to $S$. $\psi_{t}$ is $\Theta$-equivariant for all $t$. For $(x,y)\in\mathcal{W}_{1}$,
@@
As in the proof of Theorem 1.10(i) we may assume that $\Phi=\Theta$ when restricted to the Lagrangian sphere $S:x^{2}=1,y=0$ in $\mathcal{W}_{1}$, and consider the exponential map $\EuFrak{e}_{S}$ over $S$ relative to a $\Phi$-invariant $\omega_{1}$-compatible Riemannian metric $g$ on $\mathcal{W}_{1}$:
@@
$\Phi_{\theta}=\EuFrak{e}_{S}\circ\Theta_{\theta}\circ(\EuFrak{e}_{S})^{-1},\quad\forall\theta.$
@@
$F_{r}^{-1}\circ\Phi_{\theta}\circ F_{r}=\Theta_{\theta}\quad\text{on $U_{r}:=\{(x,y)\in\mathcal{W}_{1}|\ |y|^{2}\leq r\}$}.$
By allowing $r_{1}$ and $r_{0}$ to be arbitrarily large we conclude that any semi-free special Hamiltonian $S^{1}$-action on $(\mathcal{W}_{1},\omega_{1})$ is linearly up to conjugation and symplectically equivariant to the standard linear action of $\Theta$ on $U_{r}$ for any prescribed $r>0$. This completes the proof of Theorem 1.10. $\Box$
@@ 6 Non-simply connected cases
Now we consider the case that $(M,\omega)$ is a connected exact symplectic 4-manifold with $c_{1}(M,\omega)=0$ but not 1-connected. Assume that $(M,\omega)$ equips with a semi-free special Hamiltonian $S^{1}$-action which we denote as $\Phi\subset\text{Ham}(M,\omega)$.
Since the Maslov condition applies only when nonconstant orbits of $\Phi$ are homologically trivial, we assume that $Fix(\Phi)$ is nonempty and finite to accommodate the Maslov condition . Let
$k=n+1:=|Fix(\Phi)|\geq 1$
denote the number of fixed points of the $\Phi$-action, $h:M\to\mathbb{R}$ the moment map associated to $\Phi$, and $X=X_{h}$ the Hamiltonian vector field of $h$ defined by $\omega(X,\cdot)=-dh$.
Fix any $\lambda\in\Omega^{1}(M)$ with $d\lambda=\omega$. As in the 1-connected case we may assume that $\lambda$ is $\Phi$-invariant and $Fix(\Phi)\subset h^{-1}(0)$, then $\lambda$ satisfies $\lambda(X)=h$. recall
$\EuFrak{r}_{c}:=h^{-1}(c)/\Phi,\quad\text{ also \ let}\quad\EuFrak{r}^{-1}_{[-c,c]}:=h^{-1}([-c,c])$
associated to the $\Phi$-action on $M$. $h^{-1}(c)/\Phi$ is topologically an oriented surface of genus $g$ with $k$ marked points (which are the fixed points of the $\Phi$-action) and $b\geq 1$ connected boundary components at infinity. Since $\Phi$ acts freely on $h^{-1}(c)$ for all $c\neq 0$, $\EuFrak{r}_{c}=h^{-1}(c)/\Phi$ is not 1-connected, hence
@@
Recall that $\EuFrak{r}_{c}=h^{-1}(c)/\Phi$ and and $\EuFrak{r}_{c^{\prime}}=h^{-1}(c^{\prime})/\Phi$ are homeomorphic topological surfaces for $c,c^{\prime}\in\mathbb{R}$. Let $B_{0}:=\EuFrak{r}_{0}$. Up to homotopy equivalence we may identify $M$ with the corresponding $\Phi$-invariant subdomain $B=\EuFrak{r}^{-1}_{[-c,c]}B_{0}$ for sone $c>0$, $B$ is homotopic to $M$.
Part 1: Handle decomposition of $B$ and Homology of $M$. For each of the $k=1+n$ fixed points $p_{i}$, $0\leq i\leq n$, of the $\Phi$-action on $M$, denote by $q_{i}:=\EuFrak{r}_{0}(p_{i})\in B_{0}$ the image of $p_{i}$ under the map $\EuFrak{r}_{0}:M_{0}\to M_{0}/\Phi=B_{0}$. In $B_{0}$ pick $n+1$ open discs $D_{i}$ centered at $q_{i}$, $i=0,1,\dots,n$, such that the closures of $D_{i}$’s are pairwise disjoint closed discs in $B_{0}$. As $B_{0}$ can be constructed by attaching 2-dimensional 1-handles to 2-dimensional 0-handles, we may view $D_{i}$ as the set of 2-dimensional 0-handles of $B_{0}$, and with $q_{i}$ as the core point of $D_{i}$. Then $B_{0}$ can be obtained by attaching the following two types of 1-handles to $\cup_{i}D_{i}$:
@@
    Let $\tau_{i}\subset B_{0}\setminus D_{0}$, $i=1,\dots,n$, be a set of pairwise disjoint embedded arcs with endpoints such that $\tau_{i}$ is disjoint from $q_{l}$ if $l\neq i$, $q_{i}=\partial_{+}\tau_{i}$ is one endpoint of $\tau_{i}$, and $\tau_{i}$ intersects transversally with $\partial D_{0}$ at the other endpoint $\partial_{-}\tau_{i}$ of $\tau_{i}$. Thicken each of $\tau_{i}$ a bit to get $n$ mutually disjoint 2-dimensional 1-handles $H_{\tau_{i}}$ with $\tau_{i}$ as the core curve of $H_{\tau_{i}}$, $H_{\tau_{i}}\cap D_{l}=\emptyset$ if $l\neq i$ or 0, and $H_{\tau_{i}}\cap D_{0}$ is the end interval of $H_{\tau_{i}}$ which contains the endpoint $\partial_{-}\tau_{i}$ of $\tau_{i}$. WLOG we may assume that that $D_{i}\cup H_{\tau_{i}}$ is 1-connected for $i=1,2,\dots,n$.
@@
    $D$ is diffeomorphic to a 2-disc. Note that $Fix(\Phi)\subset\EuFrak{r}^{-1}(D)$, and $\EuFrak{r}^{-1}_{[-c,c]}(D)$ can be identified with a 1-connected Stein domain $W_{n}\subset\mathcal{W}_{n}$.
@@
    A disjoint union (2g+b1)(2g+b-1) 1-handles with both ends attached to $\partial D_{0}$, missing all of the handles $H_{\tau_{i}}$ in (i):
    Denote these 1-handles as $H_{C_{j}}$ for $j=1,2,\dots,2g+b-1$, where $C_{j}$ is the core arc of $H_{C_{j}}$. Denote the boundary of $H_{C_{j}}$ as $\partial H_{C_{j}}=\partial_{+}H_{C_{j}}\cup\partial_{-}H_{C_{j}}$, $H_{C_{j}}$ is attached to $\partial D$ along $\partial_{\pm}H_{C_{j}}$, where $\partial_{\pm}H_{C_{j}}$ is the disjoint union of two short intervals containing $\partial C_{j}$. By handle sliding along $\partial D$ (and thinning $\tilde{H}_{\tau_{i}}$ and $H_{C_{j}}$ if necessary) we may assume that all $\partial H_{C_{j}}$ are attached to $\partial D_{0}$ and not touching any part of $\partial H_{\tau_{i}}$.
@@
Recall $\EuFrak{r}_{0}:h^{-1}(0)\to h^{-1}(0)/\Phi\cong B_{0}$ the standard projection. Fix a $\Phi$-invariant $\omega$-compatible Riemannian metric on $M$ and let $\nabla h$ denote the gradient vector field of the moment map $h$ with respect to the metric. $M$ deformation retracts to the hypersurface $h^{-1}(0)$ via flows of gradients $\pm\nabla h$ of $h$, therefore $H_{i}(M,\mathbb{Z})\cong H_{i}(h^{-1}(0),\mathbb{Z})$ for all $i$.
Now $h^{-1}(0)=\EuFrak{r}_{0}^{-1}(B_{0})$. As $\Phi$ acts freely on $\EuFrak{r}_{0}^{-1}(B_{0})\setminus Fix(\Phi)$, we have the following results.
@@
    $\EuFrak{r}_{0}^{-1}(\tau_{i})\subset h^{-1}(0)$ is Lagrangian disc centered at $p_{i}$, $1\leq i\leq n$. If we extend $\tau_{i}$ to an embedded arc $\tilde{\tau}_{i}\subset D$ with $q_{i}$ and $q_{0}$ as endpoints, then $\tilde{S}_{i}:=\EuFrak{r}_{0}^{-1}(\tilde{\tau}_{i})$ is a Lagrangian 2-sphere with points $p_{i}$ and $p_{0}$ as poles for $i=1,2,\dots n$. These $n$ spheres are homologically independent in $\EuFrak{r}^{-1}_{0}(D)\subset h^{-1}(0)$. Indeed $\EuFrak{r}^{-1}_{0}(D)$ is homotopic to the bouquet $\cup_{i=1}^{n}\tilde{S}_{i}$ of $n$ 2-spheres.
@@
    $\EuFrak{r}_{0}^{-1}(C_{j})\subset h^{-1}(0)$ is Lagrangian cylinder with boundary attached to $\EuFrak{r}^{-1}_{0}(\partial D)$. If we extend $C_{j}$ to a simple closed curve $\tilde{C}_{j}\subset D\setminus\{q_{i}\mid i=0,\dots n\}$ then we get an embedded Lagrangian torus $\tilde{T}_{j}\subset h^{-1}(0)=\EuFrak{r}^{-1}_{0}(B_{0})$ for $j=1,2,\dots,2g+b-1$. These curves $\tilde{C}_{j}$, $j=1,\dots,2g+b-1$, together represent a basis of the homology group $H_{1}(h^{-1}(0))=H_{1}(B,\mathbb{Z})\cong\mathbb{Z}^{2g+b-1}$. As $\Phi$-orbits in $h^{-1}(0)$ contracts to points in $Fix(\Phi)$, $H_{1}(h^{-1}(0),\mathbb{Z})\cong H_{1}(B,\mathbb{Z})$ and is generated by a lifting of the curves $\tilde{C}_{j}$ in $\tilde{T_{j}}$, $j=1,\dots,2g+b-1$. The tori $\tilde{T}_{j}$’s are also independent on $H_{2}(h^{-1}(0),\mathbb{Z})$. Also $\tilde{T_{j}}$’s and $\tilde{S}_{i}$’s are pairwise linearly independent in $H_{2}(h^{-1}(0),\mathbb{Z})$, and forming a basis for $H_{2}(h^{-1}(0),\mathbb{Z})\cong\mathbb{Z}^{2g+b-1+n}$. Therefore
[…]
    In particular the homology groups of $(M,\omega,\Phi)$ is completely determined by the triple $(k=n+1,g,b)$ with $k\geq 0$, $g\geq 0$, $b\geq 1$, and $g+b\geq 2$.
Parr 2: Stein structure on $(M,\omega=d\lambda,\Phi)$
Below we study the existence of Stein structure on $(M,\omega=d\lambda,\Phi)$. Recall the following theorem due to Eliashberg [2] (see also [3] Theorem 1.3 and [4] Theorem 2.3) about the existence of a Stein structure on a 4-manifold via handle attaching criteria, which can be stated as the following after Gompf ([4] Theorem 2.3):
@@
Below we will show that the handle decomposition of $B_{0}=D_{0}\cup(\cup_{i=1}^{n}\tilde{H}_{\tau_{i}})\cup(\cup_{j}H_{C_{j}})$ lifts to a corresponding handle decomposition of $M$ as a union of handles
@@
where $\textsf{H}_{0}=\EuFrak{r}^{-1}(D_{0})$ is a 0-handle, $\textsf{H}_{V_{i}}$ are 2-handles associated to $\tilde{H}_{\tau_{i}}$, and for each $j$, the pair $(\textsf{H}_{\gamma_{j}},\textsf{H}_{U_{j}})$, where $\textsf{H}_{\gamma_{j}}$ is a 1-handle and $\textsf{H}_{U_{j}}$ is a 2-handle, is associated to $H_{C_{j}}$. Moreover, all the 2-handles $\textsf{H}_{V_{i}}$ and $\textsf{H}_{U_{j}}$ of $M$ satisfy the contact framing condition, hence $M$ admits the structure of a Stein manifold.
Case 1: $\textsf{H}_{V_{i}}$. Denote by $V_{i}:=\EuFrak{r}^{-1}_{0}(\tau_{i})$ the Lagrangian disc in $h^{-1}(0)$ with the boundary circle $\partial V_{i}$ attached to $\EuFrak{r}^{-1}_{0}(\partial D_{0})$. $V_{i}$ is the core disc of a 4-dimensional handle $\textsf{H}_{V_{i}}$ which can be identified with a subdomain of the cotangent bundle over $V_{i}$. We may assume that these handles are mutually disjoint when attached to $\EuFrak{r}^{-1}(\partial D_{0})$.
Pick any $V_{i}$ and denote it as $V$, also denote $\tau:=\tau_{i}=\EuFrak{r}(V)$, $p:=p_{i}$. We may parametrize $V$ as
$V=\{(x_{1},x_{2})\mid x_{1}^{2}+x_{2}^{2}\leq 1\},\quad p=(0,0)\in Fix(\Phi)$
@@
$\textsf{H}_{V}:=\EuFrak{r}^{-1}_{0}(H_{\tau})\cong V\times D^{2}_{y}\subset T^{*}L,\quad D^{2}_{y}=\{(y_{1},y_{2})\mid y_{1}^{2}+y_{2}^{2}<\epsilon^{2}\}$
@@
Identify $\textsf{H}_{V}$ as a subdomain of $\mathbb{C}^{2}$. Then on $\textsf{H}_{V}$ we may take $\omega=d\lambda$, with $\lambda=\frac{1}{2}\sum_{i=1}^{2}(x_{i}dy_{i}-y_{i}dx_{i})$, identify the $\Phi$-action with the standard $\Theta$-action on $\mathbb{C}^{2}$, and take the Euclidean metric as the $\Phi$-invariant $\omega$-compatible Riemannian metric on $\textsf{H}_{V}$. Then
@@
Note that both $X$ and $\nabla h$ are tangent to $\EuFrak{r}^{-1}(\partial D_{0})$ and hence to $\partial_{-}\textsf{H}_{V}$. The 1-form $\lambda$ restricts to a contact 1-form $\alpha$ on $\partial V\times D^{2}_{\epsilon}$. $\lambda(X)=0$ along $\partial V$ so $\partial V$ is a Legendrian curve to the contact form. As the outward normal vector field to $\EuFrak{r}^{-1}(D_{0})=W_{0}$ near $\partial V$ is $x_{1}\partial_{x_{1}}-x_{2}\partial_{x_{2}}\subset\ker\lambda|_{\partial V}$, the contact plane field along $\partial V$ is spanned by the symplectic pair $\{\nabla h,X\}$. So the contact framing along $\partial V$ (oriented by $X$) is given by the vector field $\nabla h|_{\partial V}=-x_{1}\partial_{y_{2}}+x_{2}\partial_{y_{1}}$.